OpenStudy (anonymous):

Let u = <-6, -2>, v = <-2, 3>. Find -3u + 2v

OpenStudy (amistre64):

multiply, and add ... seems basic enough

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ \square}} <a,b>+{\color{brown}{ \square}} <c,d>\implies <{\color{blue}{ \square}}\cdot a,{\color{blue}{ \square}}\cdot b>+ <{\color{brown}{ \square}}\cdot c,{\color{brown}{ \square}}d> \\ \quad \\ <{\color{blue}{ \square}}\cdot a+{\color{brown}{ \square}}\cdot c , {\color{blue}{ \square}}\cdot b+{\color{brown}{ \square}}d>\)

OpenStudy (amistre64):

u = <-6, -2> can be written as -6x-2y v = <-2, 3> can be written as -2x+3y what do we get with: -3(-6x-2y) + 2(-2x+3y)

OpenStudy (anonymous):

@amistre64 so we plug the other x and y values into the equation

OpenStudy (amistre64):

no, the xy parts are simply visual cues, they could be called ij or rs or whatever, but the simply tell us that the coefficient is an x component or a y componet

OpenStudy (amistre64):

we simply need to expand, collect like terms, and see what the coefficients turn out to be

OpenStudy (anonymous):

so you get 14 ?

OpenStudy (amistre64):

-3(-6x-2y) + 2(-2x+3y) 18x+6y -4x +6y 18x - 4x +6y +6y (18- 4)x +(6+6)y 14x +12y putting it back in component form we get: (14,12)

OpenStudy (anonymous):

oh okay i see now, thank you so much @amistre64

OpenStudy (amistre64):

good luck ;)